Integral x power 6 + 1\x power 2 + 1 dx please help me?

Grade:12

1 Answers

RAMESH KUMAR
6 years ago
\int\left ( x^{6}+1/x^{2}+1 \right )dx = x^{7}/7-1/x+x This is very easy, dear... Plz Understand basic of Integrations... Then also, If you have any problem, you can take guidance through Askiitians....